# WEST: LLM based Speech Toolkit for Speech Understanding, Generation, and Interaction

### v1

@@ 3.2 Speech Understanding and Question Answering
As speech understanding, we train a question answering task on Belle-1.4M-SLAM-Omni[40]. The training is divided into two stages. In the first stage, we train on AIShell-2[41] for speech recognition. In the second stage, we train on Belle-1.4M-SLAM-Omni for question answering. We also add AIShell-2 data in the second stage to avoid the loss of speech recognition ability. We apply the firered model as the speech encoder and Qwen3-1.7B as the LLM, and the model is trained on 4 A800 GPUS with pack size 25000 and 20000 steps.
We also design a Chinese question answering test set with 300 samples, which covers traditional Chinese culture, geography, STEM(Science, Technology, Engineering, Mathematics) knowledge. We name it Chinese QA test set, which is publicly available at https://github.com/wenet-e2e/west/(TODO: Fix the url). We evaluate the question answering ability of models on this test set.

### v2

@@ 3.2 Speech Understanding and Question Answering
As speech understanding, we train a question answering task on Belle-1.4M-SLAM-Omni[40]. The training is divided into two stages. In the first stage, we train on AIShell-2[41] for speech recognition. In the second stage, we train on Belle-1.4M-SLAM-Omni for question answering. We also add AIShell-2 data in the second stage to avoid the loss of speech recognition ability. We apply the firered model as the speech encoder and Qwen3-1.7B as the LLM, and the model is trained on 4 A800 GPUS with pack size 25000 and 20000 steps. During the testing stage, we use beam search decoding. For detailed configuration, please refer to https://github.com/wenet-e2e/west/blob/main/examples/belle_1.4M_qa/conf/generation_config.json.
We also design a Chinese question answering test set with 300 samples, which covers traditional Chinese culture, geography, STEM(Science, Technology, Engineering, Mathematics) knowledge. We name it Chinese QA test set, which is publicly available at https://huggingface.co/datasets/wenet-e2e/chinese_qa. We evaluate the question answering ability of models on this test set.
@@ 3.4 Speech Chat
We trained TouchChat model using the AIShell-2, Wenetspeech4TTS and Belle-1.4M-SLAM-Omni datasets. The training is divided into three stages. In the first stage, we train TouchASU model using AIshell-2 and Belle-1.4M-SLAM-Omni datasets. In the second stage, we train TouchTTS model using Wenetspeech4TTS dataset. In the third stage, we train TouchChat using Belle-1.4M-SLAM-omni, which TouchASU is freezed. The token2wav module is based on CosyVoice-300M-SFT  [^1] https://www.modelscope.cn/models/iic/CosyVoice-300M-SFT/, which converts the discrete speech tokens generated by the TouchChat into continuous audio waveforms. During the testing phase, the thinker module uses greedy decoding, while the talker module uses sampling decoding. We evaluate the spoken question answering capability of TouchChat on Chinese QA test set. Two evaluation methods are employed: S2T, where the text responses generated by the model are evaluated directly, ans S2S, where the model’s speech responses are transcribed using Paraformer  [^2] https://www.modelscope.cn/models/iic/speech_paraformer-large_asr_nat-zh-cn-16k-common-vocab8358-tensorflow1 before evaluation.
Table 4 show the results on speech chat on WEST.
Table 4: Speech Chat Results on WEST
Dataset Thinker Talker Chinese QA test
S2T(ACC) S2S(ACC)
Belle_1.4M-SLAM-Omni Qwen3-1.7B Qwen2-0.5B 64.66% 54.33%